OpenStudy (anonymous):

1. Evaluate the expression below at x = 4. (2x - 5) (x-1)

OpenStudy (mathmale):

Simply replace each occurrence of x with the value 4: (2*4 - 5)(4 - 1) = ? Better yet, write this expression with a label: f(x) = (2x - 5)(x - 1). Then evaluate f(4) = (2*4 - 5) (4 - 1).

OpenStudy (anonymous):

9?

OpenStudy (mathmale):

Yes. f(4) = (8 - 5) (4 -1) = (3) (3) = 9.

OpenStudy (anonymous):

thanks